# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v1

@@ 4 Out-of-sample performance
Next, we establish WaDiRo-SCNN out-of-sample performance. By first computing the Rademacher complexity [40], it is then possible to bound the expected out-of-sample error of our model. We make the following additional assumptions similarly to [36].
@@
The norm of modified samples from $\mathbb{F}^{\mathcal{S}}$ are bounded above: $\|{\hat{\mathbf{z}}}\|\leq R<+\infty$.
@@
As shown previously, the SCNN training problem is similar to a linear regression. We can follow the procedure used in [36] for linear regressions to bound the performance of WaDiRo-SCNNs.
@@
Under Assumptions 1, 2, and 3, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\mathbb{F}^{*}|\mathcal{S}$ is
$\mathbb{E}^{\mathbb{F}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2{B}_{*}R}{\sqrt{N}}+B_{*}R\sqrt{\frac{8\ln(2/\delta)}{N}},$ (WaDiRo-OSP)
for a modified sample $\hat{\mathbf{z}}$ from the true distribution, that was not available for training, where ${\bm{\beta}}^{\star}$ are the optimal weights obtained from (WDR1-SCNNT) or (WDR2-SCNNT).
Moreover, for any $\zeta>\frac{2{B}_{*}R}{\sqrt{N}}+{B}_{*}R\sqrt{\frac{8\ln(2/\delta)}{N}}$, we have:
[…]
@@
=$\mathcal{H}=\left\{(\mathbf{x},y)\mapsto h_{\bm{\beta}}(\mathbf{x},y):h_{\bm{\beta}}(\mathbf{x},y)=\left|y-\mathbf{x}^{\top}\bm{\beta}\right|\right\}.$
@@
Thus, the bounds follow from [36, Theorem 4.3.3] and [40, Theorem 8]. $\square$
5 Physical constraints
@@
Compared to traditional physics-informed neural networks [15], as the training problem is a convex program, knowledge of the system can be introduced as hard, convex constraints in the training phase using the decoupled formulation of the SCNN training. Let $\hat{y}_{j}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega})\right)^{\top}\hat{\mathbf{x}}_{j},\ j\in\llbracket N\rrbracket$. We formulate the physics-constrained SCNN (PCSCNN) training problem as:
$\min_{\bm{\nu},\bm{\omega}}$ $\sum_{j=1}^{N}\left|{\bm{\beta}}^{\top}{\hat{\mathbf{z}}}_{j}\right|$ (PCSCNNT)
@@
where $\mathcal{Y}_{j}\subseteq\mathbb{R}$ is the convex intersection of all physical constraints for sample $j\in\llbracket N\rrbracket$. The constraints are guaranteed to be respected during training, which is generally not the case in common physics-informed NNs. We illustrate some convex constraints with examples:
@@
$\underline{\xi}_{j}\leq\sum_{i=1}^{j}\bm{\eta}_{i}^{\top}{\mathbf{x}}_{i}+\sum_{i=1}^{j}{\gamma}_{i}\hat{y}_{i}\leq\bar{\xi}_{j}\quad$ $\forall j\in\llbracket N\rrbracket$ (7)

### v2

@@ 4 Out-of-sample performance
Next, we obtain formal out-of-sample performance guarantees for WaDiRo-SCNNs. By first computing the Rademacher complexity (Bartlett and Mendelson 2002), it is then possible to bound the expected out-of-sample error of our model. We make the following additional assumptions similarly to Chen and Paschalidis (2020).
@@
The norm of modified samples from $\mathbb{F}^{\mathcal{S}}$ are bounded above: $\|{\hat{\mathbf{z}}}\|\leq\hat{R}<+\infty$.
@@
As shown previously, the SCNN training problem is similar to a linear regression. We can follow the procedure used in Chen and Paschalidis (2020) for linear regressions to bound the performance of WaDiRo-SCNNs.
@@
Under Assumptions 1, 2, and 3, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\mathbb{F}^{\mathcal{S}}$ is:
$\mathbb{E}^{\mathbb{F}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2{B}_{*}\hat{R}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}},$
for a modified sample $\hat{\mathbf{z}}$ from the true distribution, that was not available for training, where ${\bm{\beta}}^{\star}$ are trained weights obtained from either (WDR1-SCNNT) or (WDR2-SCNNT).
Moreover, for any $\zeta>\frac{2{B}_{*}\hat{R}}{\sqrt{N}}+{B}_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}}$, we have:
$\mathrm{Pr}\left(\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\geq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\zeta\middle|\mathcal{S}\right)\leq\frac{\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2{B}_{*}\hat{R}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}}}{\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\zeta}.$
@@
=$\mathcal{H}=\left\{(\mathbf{x},y)\to h_{\bm{\beta}}(\mathbf{x},y):h_{\bm{\beta}}(\mathbf{x},y)=\left|y-\mathbf{x}^{\top}\bm{\beta}\right|\right\}.$
@@
Thus, the bounds follow from Theorem 4.3.3 of Chen and Paschalidis (2020). $\square$
Interestingly, we can obtain similar performance bounds with respect to the original distribution $\mathbb{P}$ instead of the modified one. We now make further assumptions.
Assumption 4.
The norm of the original samples, features, and labels are bounded above: $\|\mathbf{z}\|\leq R<+\infty,\ \|\mathbf{x}\|\leq S<+\infty,\text{and }\|\mathbf{y}\|\leq T<+\infty$.
Consider the general family of loss functions for arbitrary weights $\bm{\beta}$:
$\mathcal{G}=\left\{(\mathbf{x},y)\mapsto g_{\bm{\beta}}(\mathbf{x},y):g_{\bm{\beta}}(\mathbf{x},y)=\left|y-{\bm{\beta}}^{\top}\hat{\mathbf{x}}\right|,\ \hat{\mathbf{x}}=\left(\operatorname{vec}(\mathbf{x}\mathbf{d}^{\top})^{\top},\operatorname{vec}(\mathbf{x}\mathbf{d}^{\top})^{\top}\right)\right\},$ (4)
where $\mathbf{d}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$ and $\mathbf{s}_{i}\in\mathcal{S}\ \forall i\in\llbracket P\rrbracket$ are general sampling vectors.
Consider $\psi_{1}=B_{*}2PR$ for (WDR1-SCNNT) and $\psi_{2}=B_{*}\sqrt{2P}R$ for (WDR2-SCNNT), upper bounds on the prediction error of the WaDiRo-SCNN for any sample from the sample space. Then, we can rewrite the out-of-sample performance of trained WaDiRo-SCNNs as follows.
Proposition 3.
Under Assumptions 1, 3, and 4, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\mathbb{P}^{\mathcal{S}}$ is:
$\mathbb{E}^{\mathbb{P}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2\psi_{l}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}},$ (5)
for a modified sample $\hat{\mathbf{z}}$ from the true distribution given $\mathcal{S}$, that was not available for training, where ${\bm{\beta}}^{\star}$ are the optimal weights obtained from either (WDR1-SCNNT) or (WDR2-SCNNT), $\mathbb{P}^{\mathcal{S}}$ is the conditioned underlying distribution, and where $l\in\{1,2\}$.
Proof: We first bound the value of the loss function for a single arbitrary sample. For (WDR2-SCNNT) the dual norm of the weights is given by $\|\cdot\|_{2}$, thus:
$|\bm{\beta}^{\top}\hat{\mathbf{z}}|$ $\leq\|\bm{\beta}\|_{2}\|\hat{\mathbf{z}}\|_{2}$ (Hölder’s inequality)
$\leq B_{*}\sqrt{\sum_{i=1}^{P}\sum_{k=1}^{d}2(\mathbf{x}_{k}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{i}\geq 0))^{2}+y^{2}}$
$\leq B_{*}\sqrt{\sum_{i=1}^{P}\sum_{k=1}^{d}2(\mathbf{x}_{k})^{2}+y^{2}}$ (Full activation)
$=B_{*}\sqrt{2P\|\mathbf{x}\|_{2}^{2}+y^{2}}$
$\leq B_{*}\sqrt{2PS^{2}+T^{2}}$
$\leq B_{*}\sqrt{2P}R$
$=\psi_{2}.$ (Definition)
For (WDR1-SCNNT), the dual norm is the $\|\cdot\|_{\infty}$-norm. Hölder’s inequality similarly yields:
$|\bm{\beta}^{\top}\hat{\mathbf{z}}|$ $\leq\|\bm{\beta}\|_{\infty}\|\hat{\mathbf{z}}\|_{1}$ (Hölder’s inequality)
$\leq B_{*}\sum_{i=1}^{P}\sum_{k=1}^{d}2|(\mathbf{x}_{k}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{i}\geq 0))|+|y|$
$\leq B_{*}(2P\|\mathbf{x}\|_{1}+|y|)$
$\leq B_{*}(2PS+T)$
$\leq B_{*}2PR$
$=\psi_{1}.$ (Definition)
We now bound the Rademacher complexity $\mathcal{R}_{N}$ over $N$ samples for the family of functions $\mathcal{G}$ using the definition provided by Bartlett and Mendelson (2002):
$\mathcal{R}_{N}(\mathcal{G})$ $=\mathbb{E}\left[\mathbb{E}\left[\sup_{g\in\mathcal{G}}\frac{2}{N}\left|\sum_{j=1}^{N}\sigma_{j}g_{\beta}(\mathbf{x}_{j},y_{j})\right|\big{|}(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{N},y_{N})\right]\right]$
$\leq\mathbb{E}\left[\frac{2}{N}\left|\sum_{j=1}^{N}\sigma_{j}\psi_{l}\right|\right]$ (Upper bound)
$=\frac{2\psi_{l}}{N}\mathbb{E}\left[\left|\sum_{j=1}^{N}\sigma_{j}\right|\right]$
$\leq\frac{2\psi_{l}}{N}\mathbb{E}\left[\sqrt{\sum_{j=1}^{N}\sigma_{j}^{2}}\right]$ ($\sqrt{N}\geq 0$)
$=\frac{2\psi_{l}}{\sqrt{N}},$
where $\sigma_{j}\ j\in\llbracket N\rrbracket$ are independent random variables sampled from the set $\{-1,1\}$, and $l\in\{1,2\}$ accordingly to the training procedure.
Thus, by adapting Theorem 8 of Bartlett and Mendelson (2002) to loss functions mapping between $[0,B_{*}\hat{R}]$ and using the fact that $\mathcal{R}_{N}(\mathcal{G})$ is greater than the Rademacher complexity term defined in Theorem 8, we obtain the expression from (5). $\square$
We remark that the expected out-of-sample error reduces to the training error as the number of training samples grows. We also remark that the tighter the bounds are on the sample space, i.e., $\hat{R},\ R,\ S,\text{ and}\ T$, the tighter the guarantees are. Proof of Proposition 3 also leads to a probabilistic guarantee similar to the one in Proposition 2.
The theoretical guarantees provided in this section contain a probabilistic term and may be difficult to leverage empirically. Nonetheless, there is value in knowing that they exist when deploying complex ML models such as ours in critical applications because of the general lack of trust towards them. They also indicate how one can increase reliability by making these bounds tighter.
5 Physics-constrained shallow convex neural networks
@@
Compared to traditional physics-informed neural networks (Cuomo et al. 2022), because the training problem is a convex program, knowledge of the system can be introduced as hard, convex constraints in the training phase. Let $\hat{y}_{j}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega})\right)^{\top}\hat{\mathbf{x}}_{j},\ j\in\llbracket N\rrbracket$ be the prediction of the SCNN for a specific modified input vector. We formulate the physics-constrained SCNN (PCSCNN) training problem as:
$\min_{\bm{\nu},\bm{\omega}}$ $\sum_{j=1}^{N}\mathcal{L}\left(\hat{y}_{j},y_{j}\right)$ (PCSCNNT)
@@
where $\mathcal{Y}_{j}\subseteq\mathbb{R}$ is the convex intersection of all physical constraints for sample $j\in\llbracket N\rrbracket$. The constraints are guaranteed to be respected during training, which is generally not true in common physics-informed NNs. We illustrate some convex constraints with examples:
@@
$\underline{\xi}_{j}\leq\sum_{i=1}^{j}\bm{\eta}_{i}^{\top}{\mathbf{x}}_{i}+\sum_{i=1}^{j}{\gamma}_{i}\hat{y}_{i}\leq\bar{\xi}_{j}\quad$ $\forall j\in\llbracket N\rrbracket,$ (8)